1. Three fair coins are tossed. Determine the probability of each event
a) No heads
b) Exactly two head
c) Exactly one head (full answer)

Answer by greenestamps(13208)



List the possible outcomes and find the probabilities by counting the numbers of outcomes with different numbers of heads.

HHH -- 1 outcome with 3 heads
HHT
HTH
THH -- 3 outcomes with 2 heads
HTT
THT
TTH -- 3 outcomes with 1 head
TTT -- 1 outcome with 0 heads

Total possible outcomes: 2^3 = 1+3+3+1 = 8

ANSWERS:
a) 1/8
b) 3/8
c) 3/8